OpenStudy (anonymous):

WILL GIVE A MEDAL PLEASE HELP. Give the values of a, b, and c needed to write the equation's general form. 2/3(X-4)(X+5)=1 A. A = 2/3; B = 1; C = -20 B. A = 2; B = 2; C = 43 C. A = 2; B = 2; C = -43

OpenStudy (aaliyahfairgood):

ill try to help

OpenStudy (aaliyahfairgood):

2/3* (x^2+5x-4x-20)=1 2/3x+10/3-8/3-40/3=1 2/3x+2/3-40=1

OpenStudy (danjs):

I am thinking they are asking: General Form of: Ax^2 + Bx + C = 0

OpenStudy (danjs):

The way you typed it: \[\frac{ 2 }{ 3 } * (x-4)(x+5) = 1\]

OpenStudy (fibonaccichick666):

I concur there, and this may be incorrectly typed above?

OpenStudy (fibonaccichick666):

if not

OpenStudy (danjs):

\[\frac{ 2 }{ 3 } (x^2 + x - 20) = 1\] \[\frac{ 2 }{ 3 } x^2 + \frac{ 2 }{ 3 }x - \frac{ 40 }{ 3 } = 1\] \[\frac{ 2 }{ 3 } x^2 + \frac{ 2 }{ 3 }x - \frac{ 43 }{3 } = 0\] A = 2/3 B = 2/3 C = -40/3

OpenStudy (danjs):

If it is some other equation from mistype, just follow the same procedure.